# Restivo Salemi property for $\alpha$-power free languages with $\alpha\geq 5$ and $k\geq 3$ letters

### v1

@@ 1 Introduction
Let $r$ be a non-empty word, let $\alpha$ be a positive rational constant, and let $r^{\alpha}=rr\cdots rt$ be such that $\frac{|r^{\alpha}|}{|r|}=\alpha$ and $t$ is a prefix of $r$. We call the word $r^{\alpha}$ an $\alpha$-power of $r$. For example $(1234)^{3}=123412341234$, $(1234)^{\frac{7}{4}}=1234123$, and $(1234)^{\frac{1}{2}}=12$.
The square free ($2$-power free) and cube free words ($3$-power free) are well known examples of the power free words. More detailed introduction into the field of power free words can be found, for instance, in [2] and [8].
In 1985, Restivo and Salemi presented a list of five problems concerning the extendability of power free words [3]. Problem $4$ and Problem $5$ are relevant for the current article, hence we recall these two problems:
  • Problem $4$: Given finite $\alpha$-power-free words $u$ and $v$, decide whether there is a transition word $w$, such that $uwu$ is $\alpha$-power free.
@@
In 2019, a survey on the solution of all the five problems has been presented in [1]; specifically Problem $4$ and Problem $5$ were resolved for some binary language. In addition, in [1] the authors resolved Problem $4$ and $5$ for cube free words.
@@
(see [4, Remark $1$]) The definition of $\Upsilon$ says that: If $(k,\alpha)\in\Upsilon$ and $\alpha$ is a “number with $+$” then $k\geq 3$ and $\alpha\geq 2$. If $(k,\alpha)\in\Upsilon$ and $\alpha$ is “just” a number then $k=3$ and $\alpha>2$ or $k>3$ and $\alpha\geq 2$.
In 2020, Problem $4$ and Problem $5$ have were resolved for $\alpha$-power free languages over an alphabet with $k$ letters, where $(k,\alpha)\in\Upsilon$ [4].
@@
In 2018, Conjecture 1.3 appeared also in [6] using a “Restivo Salemi property”; it was defined that a language $L$ has the Restivo Salemi property if Conjecture 1.3 holds for the language $L$.
Let $L_{k,\alpha}$ denote an $\alpha$-power free language over an alphabet with $k$ letters, where $\alpha$ is a positive rational number and $k$ is positive integer.
In [6, Remark $41$], it was mentioned that for $2^{+}\leq\alpha\leq\frac{7}{3}$ the language $L_{2,\alpha}$ has the Restivo Salemi property (Conjecture 1.3 holds for $L_{2,\alpha}$), where $2^{+}$-power free language denotes the overlap free language. Otherwise, as to our best knowledge, Conjecture 1.3 remains open.
@@
Clearly $\widetilde{\Upsilon}\subset\Upsilon$. In the current article we prove that if $(k,\alpha)\in\widetilde{\Upsilon}$ then Conjecture 1.3 holds for the language $L_{k,\alpha}$.
In order to prove our result we apply the ideas and some results from [4] and [5]. We explain briefly the very basic idea. If $u,v$ are $\alpha$-power free words and $x$ is a letter such that $x$ is a factor of neither $u$ nor $v$, then clearly $uxv$ is square free ($2$-power free). Just note that there cannot be a factor in $uxv$ which is a square and contains $x$, because $x$ has only one occurrence in $uxv$. In [4], the author considers right and left infinite words containing a given factor and a non-recurrent letter $x$. Due to the non-recurrence of the letter $x$ and some involved observations of recurrent and non-recurrent factors, the author managed to prove the results. In 2023, the idea of right and left infinite power free words with a non-recurrent letter have been extended to bi-infinite $\alpha$-power free words over an alphabet with $k$-letters, where $(\alpha,k)\in\widetilde{\Upsilon}$, [5]. Moreover, in [5], it was predicted, that these bi-infinite power free words with a non-recurrent letter could help to prove Conjecture 1.3.
@@ 2 Preliminaries
Due to a significant overlap of notation, we copy and use the section “Preliminaries” from [5].
Let $\Sigma_{k}$ denote an alphabet with $k$ letters. Let $\Sigma_{k}^{+}$ denote the set of all nonempty finite words over $\Sigma_{k}$, let $\epsilon$ denote the empty word, let $\Sigma_{k}^{*}=\Sigma_{k}^{+}\cup\{\epsilon\}$, let $\Sigma_{k}^{\mathbb{N},R}$ denote the set of all right infinite words over $\Sigma_{k}$, let $\Sigma_{k}^{\mathbb{N},L}$ denote the set of all left infinite words over $\Sigma_{k}$, and let $\Sigma_{k}^{\mathbb{Z}}$ denote the set of all bi-infinite words over $\Sigma_{k}$.
@@
Let $\operatorname{occur}(w,t)$ denote the number of all occurrences of the nonempty factor $t\in\Sigma_{k}^{+}$ in the word $w\in\Sigma_{k}^{*}\cup\Sigma_{k}^{\infty}$. If $w\in\Sigma_{k}^{\infty}$ and $\operatorname{occur}(w,t)=\infty$, then we call $t$ a recurrent factor in $w$.
Let $F(w)$ denote the set of all finite factors of a finite or infinite word $w\in\Sigma_{k}^{*}\cup\Sigma_{k}^{\infty}$. The set $F(w)$ contains the empty word and if $w$ is finite then also $w\in F(w)$. Let $F_{r}(w)\subseteq F(w)$ denote the set of all recurrent nonempty factors of $w\in\Sigma_{k}^{\infty}$.
@@
[…]
@@
It is obvious that the reverse function preserves the power-freeness and that every factor of an $\alpha$-power free word is also $\alpha$-power free.
The next proposition is a “reformulation” of Corollary $1$ from [4] using only the notation of the current article.
@@
(reformulation of [4, Corollary $1$]) If $(k,\alpha)\in\Upsilon$, $v\in L_{k,\alpha}^{\mathbb{N},L}$, $z\in\operatorname{Suf}(v)$, $x\in F_{r}(v)\cap\Sigma_{k}$, $s\in L_{k,\alpha}^{\mathbb{N},L}$, and $x\not\in F(s)$ then there is a finite word $u\in\Sigma_{k}^{*}$ such that $z\in\operatorname{Suf}(su)$ and $su\in L_{k,\alpha}^{\mathbb{N},L}$.
The following elementary lemma was shown in [4]. For the reader’s convenience we present the lemma with the proof.
@@
(reformulation [4, Lemma $2$]) If $k\geq 3$ and $\alpha>2$ then $L_{k-1,\alpha}^{\mathbb{N},R}\not=\emptyset$.
Proof.
Thue Morse words are well known overlap-free ($2^{+}$-power free) right infinite words on two letters [6]. Let $x\in\Sigma_{k}$ and let $t\in L_{k,\alpha}^{\mathbb{N},R}$ be a Thue Morse word on two letters such that $x\not\in F(t)$. Since $k\geq 3$, we have that such $t$ exists. This ends the proof. ∎
Given $t\in\Sigma_{k}^{\mathbb{N},R}$, let
$\Phi(t)=\{s\in\Sigma_{k}^{\mathbb{N},L}\mid F(s)\subseteq F_{r}(t)\}\mbox{.}$
It means that all factors of $s\in\Phi(t)$ are recurrent factors of $t$. It was shown in [4], that $\Phi(t)$ is nonempty:
@@
[4, Lemma $4$] If $t\in\Sigma_{k}^{\mathbb{N},R}$ then $\Phi(t)\not=\emptyset$.
@@
If $\alpha\geq 5$, $k\geq 3$, $v\in L_{k,\alpha}^{\mathbb{Z}}$, $w\in F(v)\setminus\{\epsilon\}$, then there are $\widetilde{v}\in L_{k,\alpha}^{\mathbb{Z}}$ and a letter $x\in\Sigma$ such that $w\in F(\widetilde{v})$, $x\in F(w)$, and $\operatorname{occur}(w,x)<\infty$.
@@
However, we will not apply Theorem 2.5 directly in the current article, hence we will require some additional properties of the prefix and the suffix of the constructed bi-infinite word. Instead we will apply [5, Theorem $1$] (see Theorem 4.1 in the current article).
@@ 3 Bi-infinite $\alpha$-power words
For the rest of the article suppose that $(k,\alpha)\in\widetilde{\Upsilon}$; it means $k\geq 3$ and $\alpha\geq 5$.
We define two technical sets $\Gamma$ and $\Delta$.
@@
([5, Definition $2$]) Let $\Gamma$ be a set of triples such that $(w,\eta,u)\in\Gamma$ if and only if
@@
([5, Remark $4$]) The set $\Gamma$ contains triples of finite words $w,\eta,u$ such that $w$ is nonempty, and if $u$ is shorter than $w$, then the word $\eta$ is “sufficiently” longer than $w$.
@@
([5, Definition $3$]) Let $\Delta$ be a set of $6$-tuples such that $(s,\sigma,w,\eta,x,u)\in\Delta$ if and only if
@@
4 Transition words
@@
([5, Theorem $1$]) If $(s,\sigma,w,\eta,x,\epsilon)\in\Delta$, $t\in L_{k,\alpha}^{\mathbb{N},R}$, and $x\not\in F(t)$ then there is $\widehat{\eta}\in\operatorname{Prf}(\eta)$ such that $s\sigma w\widehat{\eta}xt\in L_{k,\alpha}^{\mathbb{Z}}$.
We show how construct an element of $\Delta$ from a bi-infinite $\alpha$-power free word $v$ with a finite factor $w$.
@@
If $v\in L_{k,\alpha}^{\mathbb{Z}}$, $w\in F(v)\setminus\{\epsilon\}$, and $x\in F(w)\cap\Sigma_{k}$ then there are $\overline{s}\in L_{k,\alpha}^{\mathbb{N},L}$, $\sigma,\eta\in\Sigma_{k}^{*}$ such that $(\overline{s},\sigma,w,\eta,x,\epsilon)\in\Delta$.
@@
    There are $\overline{v}\in L_{k,\alpha}^{\mathbb{N},L}$, $\widehat{v}\in L_{k,\alpha}^{\mathbb{N},R}$ and $\eta\in\Sigma_{k}^{*}$ such that $v=\overline{v}w\eta x\widehat{v}$ and $(w,\eta,\epsilon)\in\Gamma$.
@@
    Otherwise. Then it follows from Case 1 that in this case there are $\overline{v}\in L_{k,\alpha}^{\mathbb{N},L}$, $\widehat{v}\in L_{k,\alpha}^{\mathbb{N},R}$ and $\overline{\eta}\in\Sigma_{k}^{*}$ such that $v=\overline{v}w\overline{\eta}\widehat{v}$ such that $w\not\in F(\overline{v})$, $x\not\in F(\widehat{v})$, and $(w,\overline{\eta},\epsilon)\not\in\Gamma$.
    Let $u\in\operatorname{Prf}(\widehat{v})$ be such that $|u|>|w\overline{\eta}|$. We claim that $\overline{v}w\overline{\eta}ux\in L_{k,\alpha}^{\mathbb{N},L}$. To get a contradiction, suppose that there are $r$ and $\beta>\alpha$ such that $r^{\beta}\in\operatorname{Suf}(\overline{v}w\overline{\eta}ux)$. Since $x\not\in F(u)$ and $u\not=\epsilon$ we have that $ux\in L_{k,\alpha}$ and $ux\in\operatorname{Suf}(r)$. Since $|u|>|w\overline{\eta}|$ it follows that $w\in F(rr)$. Since $\beta>\alpha\geq 4$ and $w\not\in F(\overline{v})$, we conclude that are not such $r$ and $\beta$. Set $\eta=\overline{\eta}u$. Obviously $(w,\eta,\epsilon)\in\Gamma$, because $(w,\overline{\eta},\epsilon)\in\Gamma$.
In both cases we found $\overline{v}$ and $\eta$ such that $\overline{v}w\eta x\in L_{k,\alpha}^{\mathbb{N},L}$ and $(w,\eta,\epsilon)\in\Gamma$.
  • If $x\not\in F(\overline{v})$ then let $\overline{s}=\overline{v}$. It is easy to verify that $(\overline{s},\sigma,w,\eta,x)\in\Delta$.
  • If $x\in F(\overline{v})$ then let $t$ be a right infinite $\alpha$-power free word on the alphabet $\Sigma_{k}\setminus\{x\}$. Since $(\alpha,k)\in\widetilde{\Upsilon}$, Lemma 2.3 asserts that $t$ exists. Let $s=t^{R}$. Clearly $s$ is $\alpha$-power free and $x\not\in F(s)$.
    Let $z=w\eta x$. Proposition 2.2 implies that there is a finite word $u$ such that $su\in L_{k,\alpha}^{\mathbb{N},L}$ and $z=w\eta x\in\operatorname{Suf}(su)$. Obviously there are $\overline{s}\in L_{k,\alpha}^{\mathbb{N},L}$ and $\sigma\in\Sigma_{k}^{*}$ such that $(\overline{s},\sigma,w,\eta,x)\in\Delta$. Note that since $x\in F(w)$ and $x\not\in F(\overline{s})$ Property 4 of Definition 3.3 is easy to be asserted by a proper choice of $\sigma$.
@@
Suppose two bi-infinite $\alpha$-power free words $s_{1}\widetilde{w}_{1}t_{1}$ and $s_{2}\widetilde{w}_{2}t_{2}$, where $s_{1},s_{2}$ are left infinite words, $t_{1},t_{2}$ are right infinite words, $w_{1},w_{2}$ are finite words, the factors of $s_{2}$ are recurrent factors of $t_{1}$, and there is a letter $x$ such that $x$ is not a factor of $s_{1},s_{2},t_{1},t_{2}$ and $x$ is a factor of $w_{2}$. We show that there is a finite word $w$ such that $s_{1}\widetilde{w}_{1}w\widetilde{w}_{2}t_{2}$ is a bi-infinite $\alpha$-power free word.
@@
If $\widetilde{w}_{1},\widetilde{w}_{2}\in\Sigma_{k}^{+}$, $s_{1},s_{2}\in\Sigma_{k}^{\mathbb{N},L}$, $t_{1},t_{2}\in\Sigma_{k}^{\mathbb{N},R}$, $x\in\Sigma_{k}$, $s_{1}\widetilde{w}_{1}t_{1},s_{2}\widetilde{w}_{2}t_{2}\in L_{k,\alpha}^{\mathbb{Z}}$, $F(s_{2})\subseteq F_{r}(t_{1})$, $x\not\in F(s_{1})\cup F(s_{2})\cup F(t_{1})\cup F(t_{2})$, and $x\in F(\widetilde{w}_{2})$ then there is $w\in\operatorname{Prf}(t_{1})$ such that $s_{1}\widetilde{w}_{1}w\widetilde{w}_{2}t_{2}\in L_{k,\alpha}^{\mathbb{Z}}$.
@@
Let $p\in\operatorname{Suf}(s_{2})$ be the shortest suffix such that

### v4

@@ 1 Introduction
Due to a significant overlap of notation, we copy and use the definitions and some explications from sections “Introduction” and “Preliminaries” from [6].
Let $r$ be a non-empty word, let $\alpha$ be a positive rational number, and let $r^{\alpha}=rr\cdots rt$ be such that $\frac{|r^{\alpha}|}{|r|}=\alpha$ and $t$ is a prefix of $r$. We call the word $r^{\alpha}$ an $\alpha$-power of $r$. For example $(1234)^{3}=123412341234$, $(1234)^{\frac{7}{4}}=1234123$, and $(1234)^{\frac{1}{2}}=12$.
Suppose a finite or infinite word $w$. Let
$\begin{split}\Theta(w)=\{(r,\alpha)\mid r^{\alpha}\mbox{ is a factor of }w\mbox{ and }r\mbox{ is a non-empty word and }\\ \alpha\mbox{ is a positive rational number}\}\mbox{.}\end{split}$
We say that $w$ is $\alpha$-power-free if
$\{(r,\beta)\in\Theta(w)\mid\beta\geq\alpha\}=\emptyset$
and we say that $w$ is $\alpha^{+}$-power-free if
$\{(r,\beta)\in\Theta(w)\mid\beta>\alpha\}=\emptyset\mbox{.}$
The square free ($2$-power-free) and cube free words ($3$-power-free) are well known examples of the power-free words. More detailed introduction into the field of power-free words can be found, for instance, in [3] and [9].
In 1985, Restivo and Salemi presented a list of five problems concerning the extendability of power-free words [4]. Problem $4$ and Problem $5$ are relevant for the current article, hence we recall these two problems:
  • Problem $4$: Given finite $\alpha$-power-free words $u$ and $v$, decide whether there is a transition word $w$, such that $uwu$ is $\alpha$-power-free.
@@
In 2019, a survey on the solution of all the five problems has been presented in [2]; in particular, Problem $4$ and Problem $5$ were resolved for some binary languages. In addition, in [2] the authors resolved Problem $4$ and Problem $5$ for cube free words.
@@
([5, Remark $1$]) The definition of $\Upsilon$ says that: If $(k,\alpha)\in\Upsilon$ and $\alpha$ is a number with $+$ then $k\geq 3$ and $\alpha\geq 2$. If $(k,\alpha)\in\Upsilon$ and $\alpha$ is just a number then $k=3$ and $\alpha>2$ or $k>3$ and $\alpha\geq 2$.
In 2020, Problem $4$ and Problem $5$ were resolved for $\alpha$-power-free languages over an alphabet with $k$ letters, where $(k,\alpha)\in\Upsilon$ [5].
@@
In 2018, Conjecture 1.3 appeared also in [7] using a “Restivo Salemi property”; it was defined that a language $L$ has the Restivo Salemi property if Conjecture 1.3 holds for the language $L$. (Thus Restivo-Salemi property is exactly the transition property for extendable power-free languages.)
[…]
In [7, Remark $41$], it was mentioned that for $2^{+}\leq\alpha\leq\frac{7}{3}$ the language $L_{2,\alpha}$ has the Restivo Salemi property (Conjecture 1.3 holds for $L_{2,\alpha}$). Recall that $2^{+}$-power-free language is the overlap free language. Otherwise, as to our best knowledge, Conjecture 1.3 remains open.
@@
Clearly $\widetilde{\Upsilon}\subset\Upsilon$. In the current article, we prove that if $(k,\alpha)\in\widetilde{\Upsilon}$ then Conjecture 1.3 holds for the language $L_{k,\alpha}$; see Corollary 3.10. The requirement that $\alpha\geq 5$ was necessary in the proof of Proposition $3$ in [6]. Since we use the results of [6], we require also for the current paper that $\alpha\geq 5$.
In order to prove our result, we apply some ideas and results from [5] and [6]. We briefly explain the very basic idea. If $u,v$ are square power-free words ($2$-power-free) and $x$ is a letter such that $x$ is a factor of neither $u$ nor $v$, then clearly $uxv$ is square free. Just note that there cannot be a factor in $uxv$ which is a square and contains $x$, because $x$ has only one occurrence in $uxv$.
This very basic idea was generalized to a finite number of occurrences of $x$ instead of a one single occurrence of $x$: In [5], the author constructed right and left infinite words containing a given factor and a non-recurrent letter $x$. The non-recurrence of the letter $x$ and several involved observations of recurrent and non-recurrent factors allowed the author to resolve Problem $4$ and Problem $5$ of Restivo and Salemi for languages $L_{k,\alpha}$, where $(k,\alpha)\in\Upsilon$.
In 2023, the construction of right and left infinite power-free words with a non-recurrent letter has been extended to bi-infinite $\alpha$-power-free words over an alphabet with $k$-letters, where $(k,\alpha)\in\widetilde{\Upsilon}$ [6]. In [6], it was also predicted that these bi-infinite power-free words with a non-recurrent letter could help to prove Conjecture 1.3. In the current article we apply indeed these bi-infinite $\alpha$-power-free words with a non-recurrent letter $x$ to prove Conjecture 1.3 for languages $L_{k,\alpha}$, where $(k,\alpha)\in\widetilde{\Upsilon}$.
@@ 2 Preliminaries
Let $\Sigma_{k}^{+}$ denote the set of all non-empty finite words over $\Sigma_{k}$, let $\epsilon$ denote the empty word, let $\Sigma_{k}^{*}=\Sigma_{k}^{+}\cup\{\epsilon\}$, let $\Sigma_{k}^{\mathbb{N},R}$ denote the set of all right infinite words over $\Sigma_{k}$, let $\Sigma_{k}^{\mathbb{N},L}$ denote the set of all left infinite words over $\Sigma_{k}$, and let $\Sigma_{k}^{\mathbb{Z}}$ denote the set of all bi-infinite words over $\Sigma_{k}$.
@@
Let $\operatorname{occur}(w,t)$ denote the number of all occurrences of the non-empty factor $t\in\Sigma_{k}^{+}$ in the word $w\in\Sigma_{k}^{*}\cup\Sigma_{k}^{\infty}$. If $w\in\Sigma_{k}^{\infty}$ and $\operatorname{occur}(w,t)=\infty$, then we call $t$ a recurrent factor in $w$.
Let $F(w)$ denote the set of all finite factors of a finite or infinite word $w\in\Sigma_{k}^{*}\cup\Sigma_{k}^{\infty}$. The set $F(w)$ contains the empty word and if $w$ is finite then also $w\in F(w)$. Let $F_{r}(w)\subseteq F(w)$ denote the set of all recurrent non-empty factors of $w\in\Sigma_{k}^{\infty}$.
@@
Let $L_{k,\alpha}^{\infty}\subseteq\Sigma_{k}^{\infty}$ denote the set of all infinite $\alpha$-power-free words over $\Sigma_{k}$; formally $L_{k,\alpha}^{\infty}=\{w\in\Sigma_{k}^{\infty}\mid F(w)\subseteq L_{k,\alpha}\}$. In addition we define $L_{k,\alpha}^{\mathbb{N},R}=L_{k,\alpha}^{\infty}\cap\Sigma_{k}^{\mathbb{N},R}$, $L_{k,\alpha}^{\mathbb{N},L}=L_{k,\alpha}^{\infty}\cap\Sigma_{k}^{\mathbb{N},L}$, and $L_{k,\alpha}^{\mathbb{Z}}=L_{k,\alpha}^{\infty}\cap\Sigma_{k}^{\mathbb{Z}}$; it means the sets of right infinite, left infinite, and bi-infinite $\alpha$-power-free words, respectively.
@@
It is obvious that the reverse function preserves the power-freeness and that every factor of an $\alpha$-power-free word is also $\alpha$-power-free.
The next proposition is a reformulation of Corollary $1$ from [5] using only the notation of the current article.
@@
(reformulation of [5, Corollary $1$]) If $(k,\alpha)\in\Upsilon$, $v\in L_{k,\alpha}^{\mathbb{N},L}$, $z\in\operatorname{Suf}(v)$, $x\in F_{r}(v)\cap\Sigma_{k}$, $s\in L_{k,\alpha}^{\mathbb{N},L}$, and $x\not\in F(s)$ then there is a finite word $u\in\Sigma_{k}^{*}$ such that $z\in\operatorname{Suf}(su)$ and $su\in L_{k,\alpha}^{\mathbb{N},L}$.
Remark 2.3.
([6, Remark $3$]) Proposition 2.2 says that if $z$ is a finite $\alpha$-power-free word that can be extended to a left infinite $\alpha$-power-free word having a letter $x$ as a recurrent factor and $s$ is a left infinite $\alpha$-power-free word not containing the letter $x$ as a factor, then there is a left infinite $\alpha$-power-free word containing $z$ as a suffix and having only a finite number of occurrences of $x$.
The next elementary lemma follows easily from famous theorems of Thue [10, 11].
@@
(reformulation of [5, Lemma $2$]) If $k\geq 3$ and $\alpha>2$ then $L_{k-1,\alpha}^{\mathbb{N},R}\not=\emptyset$.
Given $t\in\Sigma_{k}^{\mathbb{N},R}$, let $\Phi(t)=\{s\in\Sigma_{k}^{\mathbb{N},L}\mid F(s)\subseteq F_{r}(t)\cup\{\epsilon\}\}\mbox{.}$ It means that all non-empty factors of $s\in\Phi(t)$ are recurrent factors of $t$. The following lemma is an immediate consequence of König’s lemma [1].
@@
([5, Lemma $4$]) If $k\in\mathbb{N}$ and $t\in\Sigma_{k}^{\mathbb{N},R}$ then $\Phi(t)\not=\emptyset$.
@@
If $(k,\alpha)\in\widetilde{\Upsilon}$, $v\in L_{k,\alpha}^{\mathbb{Z}}$, $w\in F(v)\setminus\{\epsilon\}$, then there are $\widetilde{v}\in L_{k,\alpha}^{\mathbb{Z}}$ and a letter $x\in\Sigma_{k}$ such that $w\in F(\widetilde{v})$, $x\in F(w)$, and $\operatorname{occur}(w,x)<\infty$.
@@
We will not apply Theorem 2.6 directly, because we will need some additional properties of the constructed bi-infinite word. Instead, we will apply [6, Theorem $1$] (see Theorem 3.4 in the current article).
@@ 3 Bi-infinite $\alpha$-power words
For the rest of the article suppose that $(k,\alpha)\in\widetilde{\Upsilon}$; it means $k\geq 3$ and $\alpha\geq 5$. We define two technical sets $\Gamma$ and $\Delta$.
@@
([6, Definition $2$]) Let $\Gamma$ be a set of triples such that $(w,\eta,u)\in\Gamma$ if and only if
@@
([6, Remark $4$]) The set $\Gamma$ contains triples of finite words $w,\eta,u$ such that $w$ is non-empty, and if $u$ is shorter than $w$, then the word $\eta$ is “sufficiently” longer than $w$.
@@
([6, Definition $3$]) Let $\Delta$ be a set of $6$-tuples such that
$(s,\sigma,w,\eta,x,u)\in\Delta$
if and only if
@@
In [6], it was shown that if $(s,\sigma,w,\eta,x,\epsilon)\in\Delta$ and $t$ is a right infinite $\alpha$-power-free word with no occurrence of the letter $x$ then there is a bi-infinite $\alpha$-power-free word containing the factor $w$ and having only a finite number of occurrences of $x$:
@@
([6, Theorem $1$]) If $(s,\sigma,w,\eta,x,\epsilon)\in\Delta$, $t\in L_{k,\alpha}^{\mathbb{N},R}$, and $x\not\in F(t)$ then there is $\widehat{\eta}\in\operatorname{Prf}(\eta)$ such that $s\sigma w\widehat{\eta}xt\in L_{k,\alpha}^{\mathbb{Z}}$.
Definition 3.5.
([6, Definition $4$]) Suppose $v\in L_{k,\alpha}^{\mathbb{Z}}$ and $w\in F(v)\setminus\{\epsilon\}$. If there are $v_{1}\in L_{k,\alpha}^{\mathbb{N},L}$ and $v_{2}\in L_{k,\alpha}^{\mathbb{N},R}$ such that $v=v_{1}v_{2}$ and $w\in F_{r}(v_{2})$ then we say that $w$ is on-right-side recurrent in $v$.
Remark 3.6.
[6, Remark $8$]) Note in Definition 3.5 that no restriction is imposed on the recurrence of $w$ in $v_{1}$. This means that $w$ may also be recurrent in $v_{1}$.
We show how to construct an element of $\Delta$ from a bi-infinite $\alpha$-power-free word $v$ with a finite factor $w$ containing the letter $x$.
@@
If $v\in L_{k,\alpha}^{\mathbb{Z}}$, $w\in F(v)\setminus\{\epsilon\}$, and $x\in F(w)\cap\Sigma_{k}$ then there are $\overline{s}\in L_{k,\alpha}^{\mathbb{N},L}$, and $\sigma,\eta\in\Sigma_{k}^{*}$ such that $(\overline{s},\sigma,w,\eta,x,\epsilon)\in\Delta$.
@@
    If $\operatorname{occur}(v,w)=\infty$ or $x$ is on-right-side recurrent in $v$ then it is easy to see that there are $\overline{v}\in L_{k,\alpha}^{\mathbb{N},L}$, $\widehat{v}\in L_{k,\alpha}^{\mathbb{N},R}$, and $\eta\in\Sigma_{k}^{*}$ such that $v=\overline{v}w\eta x\widehat{v}$ and $(w,\eta,\epsilon)\in\Gamma$.
@@
    Otherwise; it means that $\operatorname{occur}(v,w)<\infty$ and $x$ is not on-right-side recurrent in $v$. Then there are $\overline{v}\in L_{k,\alpha}^{\mathbb{N},L}$, $\widehat{v}\in L_{k,\alpha}^{\mathbb{N},R}$ and $\overline{\eta}\in\Sigma_{k}^{*}$ such that $v=\overline{v}w\overline{\eta}\widehat{v}$,
    $x\not\in F(\widehat{v})\mbox{, }$ (1)
    and
    $\operatorname{occur}(\overline{v}w,w)=1\mbox{.}$ (2)
    Let $\eta_{0}\in\operatorname{Prf}(\widehat{v})$ be such that
    $|\eta_{0}|>\max\{|w\overline{\eta}|,(\alpha+1)\alpha^{|w|}|w|\}\mbox{.}$ (3)
    We claim that $\overline{v}w\overline{\eta}\eta_{0}x\in L_{k,\alpha}^{\mathbb{N},L}$. To get a contradiction, suppose that there are $r$ and $\beta>\alpha$ such that $r^{\beta}\in\operatorname{Suf}(\overline{v}w\overline{\eta}\eta_{0}x)$. From (1) we have that $x\not\in F(\eta_{0})$; consequently $\eta_{0}x\in L_{k,\alpha}$ and $\eta_{0}x\in\operatorname{Suf}(r)$. From (3) we have that $|\eta_{0}|>|w\overline{\eta}|$; it follows that $w\in F(rr)$. From $\beta>\alpha\geq 5$ and (2), we conclude that are not such $r$ and $\beta$. Set $\eta=\overline{\eta}\eta_{0}$. From (3) it is obvious $(w,\eta,\epsilon)\in\Gamma$.
In both cases we found $\overline{v}$ and $\eta$ such that $\overline{v}w\eta x\in L_{k,\alpha}^{\mathbb{N},L}$, and $(w,\eta,\epsilon)\in\Gamma$.
  • If $x\not\in F(\overline{v})$ then let $\overline{s}=\overline{v}$. It is obvious that $(\overline{s},\epsilon,w,\eta,x,\epsilon)\in\Delta$.
  • If $x\in F(\overline{v})$ then let $t$ be a right infinite $\alpha$-power-free word on the alphabet $\Sigma_{k}\setminus\{x\}$. Since $(k,\alpha)\in\widetilde{\Upsilon}$, Lemma 2.4 asserts that $t$ exists. Let $s=t^{R}$. Clearly $s$ is a left infinite $\alpha$-power-free word and $x\not\in F(s)$.
    Let $z=w\eta x$. Proposition 2.2 implies that there is a finite word $u$ such that $su\in L_{k,\alpha}^{\mathbb{N},L}$ and $z\in\operatorname{Suf}(su)$. Obviously there are $\overline{s}\in L_{k,\alpha}^{\mathbb{N},L}$ and $\sigma,\eta_{0}\in\Sigma_{k}^{*}$ such that $\overline{s}\sigma w\eta_{0}x=su$, $\eta\in\operatorname{Suf}(\eta_{0})$, $w\eta x\in\operatorname{Suf}(\overline{s}\sigma w\eta_{0}x)$, and $(\overline{s},\sigma,w,\eta_{0},x,\epsilon)\in\Delta$. Note that since $x\in F(w)$ and $x\not\in F(\overline{s})$ Property 4 of Definition 3.3 is easy to be asserted by a proper choice of $\sigma$ and $\eta_{0}$.
@@
Suppose two bi-infinite $\alpha$-power-free words $s_{1}\widetilde{w}_{1}t_{1}$ and $s_{2}\widetilde{w}_{2}t_{2}$, where $s_{1},s_{2}$ are left infinite words, $t_{1},t_{2}$ are right infinite words, $\widetilde{w}_{1},\widetilde{w}_{2}$ are finite words, the factors of $s_{2}$ are recurrent factors of $t_{1}$, and there is a letter $x$ such that $x$ is not a factor of $s_{1},s_{2},t_{1},t_{2}$ and $x$ is a factor of $\widetilde{w}_{2}$. We show that there is a finite word $w$ such that $s_{1}\widetilde{w}_{1}w\widetilde{w}_{2}t_{2}$ is a bi-infinite $\alpha$-power-free word.
@@
If $\widetilde{w}_{1},\widetilde{w}_{2}\in\Sigma_{k}^{+}$, $s_{1},s_{2}\in\Sigma_{k}^{\mathbb{N},L}$, $t_{1},t_{2}\in\Sigma_{k}^{\mathbb{N},R}$, $x\in\Sigma_{k}$, $s_{1}\widetilde{w}_{1}t_{1},s_{2}\widetilde{w}_{2}t_{2}\in L_{k,\alpha}^{\mathbb{Z}}$, $F(s_{2})\subseteq F_{r}(t_{1})$, $x\not\in F(s_{1})\cup F(s_{2})\cup F(t_{1})\cup F(t_{2})$, and $x\in\operatorname{Prf}(\widetilde{w}_{2})$ then there is $w\in\operatorname{Prf}(t_{1})$ such that $s_{1}\widetilde{w}_{1}w\widetilde{w}_{2}t_{2}\in L_{k,\alpha}^{\mathbb{Z}}$.
@@
Let $p\in\operatorname{Suf}(s_{2})$ be the shortest suffix of $s_{2}$ such that